# Privacy Perspectives and Practices of Chinese Smart Home Product Teams

### v1

@@ 5.4.4 Encourage privacy training and collaboration
Participants prioritized meeting legal needs but often relied heavily on specialized compliance staff, as noted in §4.1.2. This over-reliance led to limited legal awareness among broader product teams regarding their responsibilities. While many participants expressed empathy for users’ privacy needs, this alone was insufficient without clear internal guidelines and a solid grasp of legal frameworks. To address this, we recommend collaborative work among team members early in product design, supported by strong corporate policies that guide teams through the evolving privacy challenges in Chinese multi-user smart homes. Companies should also develop structured training initiatives—such as mandatory lectures/courses that address legal compliance and ethical concerns rooted in cultural contexts [67, 3]. These should be part of on-boarding learning and ongoing professional development, featuring real-life case studies that present ethical trade-offs and cultural tensions around user privacy [3]. This helps teams turn empathy into actionable strategies by aligning legal knowledge with ethical judgment. We also encourage partnerships between companies, academia, and policymakers to co-develop standardized privacy education programs that highlight both global best practices and local cultural nuances.
@@ 5.4.5 Consider privacy compliance in other markets
As discussed in §4.3.3, participants expressed the need for a stronger understanding of legal and cultural differences across international markets (e.g., EU countries) to design smart home devices that are both compliant and aligned with local privacy norms—such as individual rights versus national security priorities. Since many Chinese-developed devices are exported globally, it is essential for marketing and compliance teams to conduct workshops that inform product teams about regional privacy laws (e.g., GDPR) and cultural expectations. We recommend adopting a user-centered, culturally-adaptive privacy-by-design framework or approach that includes conducting privacy impact assessments with culturally-specific adaptations, which would be beneficial in contexts or countries where collectivist values and weaker individual privacy regulations dominate. Engaging users from different regions early—through collaboration with local stakeholders and culturally diverse usability testing—can help identify privacy and usability concerns. Customizing privacy settings for specific markets is especially important in regions where privacy regulations are still developing, such as Southeast Asia, the Middle East, and Africa [12, 11, 97, 46, 86].
@@ 5.5 Future Work
Future research should explore privacy and security issues from the perspectives of at-risk individuals in Chinese multi-user smart home environments. Participants from Chinese smart home product teams anticipated emerging AI-driven features—such as real-time voice recognition and generative LLM-based dialogue in smart speakers and cameras [60]. Although these advances may improve UX, they also raise significant concerns about passive data collection and user profiling [142, 141]. Future work should also examine smart home privacy policies and design practices more closely, particularly the growing gaps between these practices and Chinese privacy regulations (e.g., PIPL).
@@ 6 Conclusion
Our interviews with 27 Chinese smart home product team members provide new insights into privacy perspectives and practices in a non-WEIRD context. Participants demonstrated a strong commitment to legal compliance and ethical considerations within China’s strict policy environment. However, they also faced challenges in navigating the complexities of an evolving regulatory and technological landscape, including insufficient public attention to at-risk groups, gender disparities, and power dynamics between authorities, companies, and users. In addition, we examine how cultural and social norms in China influence individuals’ perceptions of privacy. Chinese smart home product teams may often regard privacy as a lower priority for users—an attitude shaped by the collectivist orientation of Chinese society. This, in turn, may lead to privacy de-prioritization in product design and development life-cycle. Finally, we recommend the development of clear guidelines for transparent data practices in smart home devices, with an emphasis on prioritizing end-users’ privacy needs ,particularly those at heightened risk. This involves navigating a delicate balance between individual privacy rights and national security considerations. Additionally, we recommend privacy-focused training and educational resources for users and for smart home product teams in China, who may operate under the assumption that the general public does not value privacy.
Appendix A Author Positionality
Since privacy can vary significantly based on cultural, social, and individual factors, we did not provide participants with a specific definition of privacy (given privacy definitions in the academic literature are predominantly influenced by WEIRD values). Instead, we explored their privacy perspectives and practices influenced by their backgrounds and experiences, given that our work focuses on a non-WEIRD country. Although participants did not explicitly mention the term “Confucianism” during the interviews, they discussed values aligned with it, such as filial piety (see §4.2.2). Recognizing these culturally ingrained themes, we integrated Confucianism into our analysis to gain deeper insights into participants’ perspectives. Instead of treating Confucianism as a discrete code, we considered it a broader thematic lens that emerged during data analysis.
Further, our positionality and interpretation of the results are influenced by our backgrounds and experiences. All authors are experienced researchers in human-centered computing, with extensive work involving both majority populations and at-risk groups. While all authors are trained researchers currently working in predominantly Western institutions, the interviewers and data analysts bring additional insights from their personal backgrounds, having been born and raised in China. This dual perspective enables them to analyze interviews and interpret data with both academic rigor and deep cultural insight. Also, one author specializes in technology privacy law and policy in both China and Western countries, including the UK and US.
Appendix B Privacy Laws
While GDPR, CCPA, and Chinese privacy laws (e.g., PIPL) share common goals, such as protecting personal data and ensuring individuals’ rights, they differ significantly in government access, enforcement mechanisms, and the emphasis on national security versus individual rights [22]. In contrast to the EU’s approach, which is motivated by safeguarding individual rights and data privacy, China’s emphasis on cyber-sovereignty is driven primarily by national security concerns [40]. GDPR highlights extensive rights such as the right to be forgotten, data portability, and explicit consent requirements, while CCPA focuses more on access and deletion rights and the option to opt out of data selling. PIPL, influenced by GDPR, grants rights to data access, correction, deletion, and portability but includes provisions for government access [125].
Similarities Differences
Personal Information Protection and Its Scope All these laws and regulations cover personal and sensitive information protection. Several legal instruments interact with each other, potentially leading to conflicts, overlap, or uncertainty in compliance and enforcement (e.g., overlap of legislative bodies [144]). PIPL-personal data protection; CSL-overall national cybersecurity; DSL-broader data security issues. POPCPI-children data protection; PIPL, DSL, CSL-critical information infrastructure operators; PCPPI-service providers handling children’s data. The limitations can differ in scope and complexity based on the specific focus and goals of each legal instrument. For instance, the CSL and the DSL deal with broader data security issues beyond the confines of personal information.
Privacy Principles and Measures All laws emphasize privacy principles (e.g., informed consent, security measures, individual rights, and purpose limitation). Challenge of striking the right balance between privacy protection and national security or public interests, particularly evident in government access to data (e.g., strict data governance on cross-boarder data transfer vs. individual sensitive data protection [40, 125]). The standards, measures, and provisions vary in their level of specificity and detail (e.g., DSL and CSL focus on activities that may infringe upon data sovereignty, while PIPL is targeted more at specific aspects of the handling of data by data processors). PIPL, CSL and DSL govern data collected in China which may pose challenges for multinational corporations in terms of global interactions and cross-border operationality across multiple jurisdictions.
Protection of Minors’ Data PIPL and POPCPI contain provisions addressing the special protection of minors’ data (under 14); CSL highlights the need to build a safe and healthy cyber environment for minors. PIPL and POPCPI require obtaining explicit and informed consent from guardians for processing children’s data. They emphasize implementing security measures and the need for transparency in data processing to protect children’s data from unauthorized access, disclosure, alteration, or destruction. DSL does not have provisions for minors; The PIPL, DSL, and CSL are legally binding, while POPCPI is not legally binding but offers guidelines for compliance. POPCPI faces enforcement challenges in verifying children’s ages online, ensuring informed consent, and across various platforms and holding non-compliant providers accountable. Additionally, transitioning from stricter to less restrictive data protection measures as children grow older can be complex and pose privacy risks.
Cross-border Data Transfers and Data Localization All introduce some form of regulation or assessment for cross-border data transfers to safeguard data privacy and national security. Many data governance laws and regulations have developed since 2021, when cross-border data transfers became a significant national concern. These laws and regulations significantly increased the administrative hurdles and operational costs for multinational corporations to ensure compliance. DSL-stricter requirements for ‘important data’ and critical information infrastructure operators (CIIOs); CSL-security assessment; PIPL-personal information can be transferred abroad if certain conditions are met. Data localization limitations might be more significant for laws like DSL and PIPL due to their impact on cross-border data flows. Limitations related to mass surveillance are more prominent in CSL due to its specific national security focus.
Challenges and Technological Advancements Emphasize the need for data processors to adapt to fast-paced technological changes while ensuring compliance with existing laws. All these laws did not specially highlight the privacy issues from AI, IoT. PIPL highlights that rapid technological advancements can lead to challenges in addressing privacy and security issues (e.g.,from facial recognition). DSL-vague definition of ‘important data’; CSL-the rapid evolution of cyber threats require continuous updates to privacy measures, increases regulatory burden for compliance; PIPL-lacks specific guidelines for emerging technologies, introducing uncertainty about compliance.
TABLE II: Similarities and differences of Chinese privacy laws and regulations – PIPL, DSL, CSL, and POPCPI.

### v2

@@ 5.4.4 Encourage privacy training and collaboration
Participants prioritized meeting legal needs, but often relied heavily on specialized compliance staff, as noted in §4.1.2. This over-reliance led to limited legal awareness among broader product teams regarding their responsibilities. While many participants expressed empathy for users’ privacy needs, this alone was insufficient without clear internal guidelines and a solid grasp of legal frameworks. To address this, we recommend collaborative work among team members early in product design, supported by strong corporate policies that guide teams through the evolving privacy challenges in Chinese multi-user smart homes. Companies should also develop structured training initiatives—such as mandatory lectures/courses that address legal compliance and ethical concerns rooted in cultural contexts [60]. These should be part of on-boarding learning and ongoing professional development, featuring real-life case studies that present ethical trade-offs and cultural tensions around user privacy, helping teams turn empathy into actionable strategies by aligning legal knowledge with ethical judgment. We also encourage partnerships between companies, academia, and policymakers to co-develop standardized educational programs that highlight both global best practices and local cultural nuances.
@@ 5.4.5 Consider privacy compliance in other markets
As discussed in §4.3.3, participants emphasized the need for a deeper understanding of legal and cultural differences across international markets (e.g., EU countries) in order to design smart home devices that are both compliant and aligned with local privacy norms. We recommend adopting a user-centered, culturally-adaptive privacy-by-design framework or approach that includes conducting privacy impact assessments with culturally-specific adaptations, which would be beneficial in contexts or countries where collectivist values and weaker individual privacy regulations dominate. Engaging users from different regions—through collaboration with local stakeholders and culturally diverse usability testing—can help identify privacy and usability concerns. Customizing privacy settings for specific markets is especially important in regions where privacy regulations are still developing, such as Southeast Asia, the Middle East, and Africa [11, 10, 84, 44, 75].
Our findings also revealed that a small number of our participants (e.g., P12) were involved in designing smart home products for overseas customers. These participants developed stronger privacy awareness through engagement with international regulations such as GDPR. This exposure shaped their global compliance strategies and influenced domestic design practices, with GDPR often serving as a benchmark for meeting or exceeding standards in other regions, such as India or Latin America. However, this transnational orientation created tensions between international legal expectations and local user norms or marketability concerns. To navigate these challenges, participants emphasized the value of internal workshops to help teams better understand and localize privacy expectations across regions.
@@ 5.5 Future Work
Future research should investigate privacy and security issues from the perspectives of at-risk individuals in Chinese multi-user smart home environments. Participants from Chinese smart home product teams anticipated the emergence of AI-driven features—such as real-time voice recognition and generative LLM-based dialogue in smart speakers and cameras [55]. While these advancements may enhance UX, they also raise significant concerns regarding passive data collection and user profiling [127, 126]. Future work should further examine smart home privacy policies and design practices, with particular attention to the widening gaps between these practices and Chinese privacy laws and regulations (e.g., PIPL).
@@ 6 Conclusion
Our interviews with 27 members of Chinese smart home product teams provide new insights into their privacy perspectives and practices in a non-WEIRD context. Participants demonstrated a strong commitment to legal compliance and ethical considerations within China’s strict legal and regulatory environment. However, they also faced challenges in navigating the complexities of an evolving regulatory and technological landscape, including limited understanding of the privacy needs of at-risk groups, gender disparities, and power imbalances among authorities, companies, and users. Moreover, our findings highlight how cultural and social norms in China shape perceptions of privacy. Smart home product teams often perceived privacy as a lower priority for users—an attitude influenced by the collectivist orientation of Chinese society—which might have contributed to the de-prioritization of privacy in product design and development.
We recommend the creation of clear guidelines for transparent data practices in smart home devices, emphasizing the protection of end-users’ privacy, particularly for those at heightened risk. Achieving this requires balancing individual privacy rights with national security considerations. In addition, we call for privacy-focused training and educational initiatives for both users and smart home product teams in China, who may operate under the assumption that the general public places limited value on privacy.
Acknowledgments
The authors would like to thank all participants from Chinese smart home product teams for taking part in this study, as well as the anonymous reviewers for their insightful feedback. We also extend our gratitude to Yixin Zou, Julia Bernd, and Adam Jenkins for their valuable comments on earlier drafts of this paper.
Appendix A Meta-Review
The following meta-review was prepared by the program committee for the 2026 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
A.1 Summary
This paper conducts a semi-structured interview study with 27 with Chinese smart home product team members, including managers, engineers, designers, legal/privacy experts, and marketers. The paper focuses on teams’ approaches and perceptions of privacy, and the challenges they face when addressing privacy. The authors find that participants emphasized compliance with Chinese data privacy laws and situate their findings within the Chinese context, including how ideas of Confucianism and filial piety play a role in how people think about and design for privacy.
A.2 Scientific Contributions
  • Independent Confirmation of Important Results with Limited Prior Research
  • Provides a Valuable Step Forward in an Established Field
A.3 Reasons for Acceptance
  1. 1.
    This paper provides independent confirmation of important results with limited prior research by investigating how Chinese smart home product teams interact with and respond to privacy regulations.
  2. 2.
    This paper provides a valuable step forward in an established field by examining privacy in a non-Western cultural context, including Confucianism and filial piety influences on privacy and design practices.
A.4 Noteworthy Concerns
  • The interview guide may be leading, with early emphasis on privacy and legal requirements potentially priming responses to later questions on roles and data practices.